OpenStudy (anonymous):

The width and length of a rectangle are consecutive intergers. The perimeter of the rectangle is 145m. What is the width and the length?

OpenStudy (anonymous):

2(x)+2(x+1)=145

OpenStudy (anonymous):

I don't think it is possible with consecutive integers with that certain perimeter.

OpenStudy (anonymous):

143/4=x 143/4=35.75

OpenStudy (anonymous):

w - 35.75 l - 36.75

OpenStudy (across):

They're not integers though :/